BRICS Notes Series, 1996

July 7, 2003

This document is also available as PostScript and DVI.

References

[NS-96-15] Abstract, PostScript, PDF, DVI.

CoFI. CASL – The CoFI Algebraic Specification Language; Tentative Design: Language Summary. December 1996. viii+34 pp.

[NS-96-14] Abstract, PostScript, PDF, DVI.

Peter D. Mosses. A Tutorial on Action Semantics. December 1996. 46 pp. Tutorial notes for FME ’94 (Formal Methods Europe, Barcelona, 1994) and FME ’96 (Formal Methods Europe, Oxford, 1996).

[NS-96-13] Olivier Danvy, editor. Proceedings of the Second ACM SIGPLAN Workshop on Continuations, CW ’97 (ENS, Paris, France, 14 January, 1997), December 1996. 166 pp.

[NS-96-12] Mandayam K. Srivas. A Combined Approach to Hardware Verification: Proof-Checking, Rewriting with Decision Procedures and Model-Checking; Part II: Articles. BRICS Autumn School on Verification. October 1996. 56 pp.

[NS-96-11] Mandayam K. Srivas. A Combined Approach to Hardware Verification: Proof-Checking, Rewriting with Decision Procedures and Model-Checking; Part I: Slides. BRICS Autumn School on Verification. October 1996. 29 pp.

[NS-96-10] Robert Pollack. What we Learn from Formal Checking; Part III: Formalization is Not Just Filling In Details. BRICS Autumn School on Verification. October 1996. iv+42 pp.

[NS-96-9] Robert Pollack. What we Learn from Formal Checking; Part II: Using Type Theory: An Introduction. BRICS Autumn School on Verification. October 1996. iv+71 pp.

[NS-96-8] Robert Pollack. What we Learn from Formal Checking; Part I: How to Believe a Machine-Checked Proof. BRICS Autumn School on Verification. October 1996. iv+19 pp.

[NS-96-7] Tom F. Melham. Some Research Issues in Higher Order Logic Theorem Proving. BRICS Autumn School on Verification. October 1996. 15 pp.

[NS-96-6] Gerard J. Holzmann. On-the-Fly Model Checking Tutorial. BRICS Autumn School on Verification. October 1996. 31 pp.

[NS-96-5] Thomas A. Henzinger. Automatic Verification of Real-Time and Hybrid Systems. BRICS Autumn School on Verification. October 1996. 28 pp.

[NS-96-4] Edmund M. Clarke, Jr. Symbolic Model Checking. BRICS Autumn School on Verification. October 1996. 55 pp.

[NS-96-3] David A. Basin. Verification Based on Monadic Logic. BRICS Autumn School on Verification. October 1996. 43 pp.

[NS-96-2] Abstract, PostScript, PDF.

Allan Cheng, Kim G. Larsen, and Mogens Nielsen, editors. Programme and Abstracts of the BRICS Autumn School on Verification (Aarhus, Denmark, October 28 – November 1, 1996), August 1996. ii+18pp.

[NS-96-1] Abstract, PostScript, PDF.

Andr Berthiaume. Quantum Computation. Mini-Course. January 1996. iv+126 pp.